SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 3,400 of these meals using 1,300 direct labor-hours. The company paid its direct labor workers a total of $16,900 for this work, or $13.00 per hour.

According to the standard cost card for this meal, it should require 0.40 direct labor-hours at a cost of $12.00 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 3,400 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 3,400 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?
Answer #1

1) Standard labor hours allowed = 3400*.40 = 1360 Hours

2) Standard labor cost allowed = 1360*12 = 16320

3) Labor spending variance = 16320-16900 = 580 U

4) Labor rate variance = (12*1300-16900) = 1300 U

Labor efficiency variance = (1360-1300)*12 = 720 F